OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is parallel to y = 3x+6 and passes through the point (-10, 2.5) Whoever awnsers gets a medal :D

OpenStudy (anonymous):

y = 3x + 6....the slope here is 3 and a parallel line will have the same slope. Now we will use point slope formula : y - y1 = m(x - x1) slope(m) = 3 (-10, 2.5) x1 = -10 and y1 = 2.5 now sub y - 2.5 = 3(x - (-10) y - 2.5 = 3(x + 10) y - 2.5 = 3x + 30 y = 3x + 30 + 2.5 y = 3x + 32.5 <--- your parallel line

OpenStudy (anonymous):

Thanks again :D

OpenStudy (anonymous):

no problem :)